OpenStudy (anonymous):

determine the constant k so that the graph of 2(y+3y)=3k(x-1) will have the slope of 6

OpenStudy (dan815):

write in y=mx+b form

OpenStudy (dan815):

2*4y=3k(x-1) how can u write this as y=something

OpenStudy (nincompoop):

nvvm

OpenStudy (anonymous):

\[ 2(y+3y)=3k(x-1) \]Same as \[ 8y=3k(x-1) \]Differentiate both sides \[ 8y' = 3k \]We want \(y'=6\).\[ 8(6) = 3k \]Solve for \(k\). I think.

OpenStudy (nincompoop):

I didn't even know you can do it that way @wio good to know and thank you

OpenStudy (anonymous):

16